OpenStudy (anonymous):

what are the zeros of y=X^2-4x-5 1,-5 1,5 -1,-5 -1,5

OpenStudy (luigi0210):

Factor the equation: \[\LARGE x^2-4x-5=(x-5)(x+1)\] Now solve this: \[\LARGE (x-5)(x+1)=0\]

OpenStudy (anonymous):

got it thanks!

OpenStudy (luigi0210):

Yup~

OpenStudy (anonymous):

D